# Through the Heliospheric Lens: Directional Deflection of High-Energy Cosmic-Ray Electrons and Positrons

### v1

@@ 4.4 HCS Geometry: Tilt and Waviness
Comparing panels 5(d) and 5(c) shows that replacing a flat sector boundary with a tilted, wavy sheet broadens the 16–84% bands—especially at $E\!\sim\!30$$300$ GeV—while leaving the mean only modestly affected. Physically, additional undulations increase the chance and multiplicity of sector crossings and drift-rich path segments; small changes in launch direction translate into larger path-length and curvature differences across the sky.
@@
Strengthening the local azimuthal field (panel 5(e)) modestly raises sub-TeV means relative to the flat-HCS DQCS+spiral case, consistent with tighter winding, but the difference is again minor compared to the Parker vs. DQCS split from Sec. 4.3. At TeV energies the mean curves nearly coincide.
@@
Introducing a second wave in longitude (panel 5(f)) further amplifies the percentile width at low energies while keeping the mean close to the $m=1$ case. The mean-overlay in Fig. 6 highlights that ordering: Parker remains largest, DQCS(core) is the most suppressed, and the DQCS+spiral variants (flat/wavy, SB) cluster together in mean but differ in spread.
@@ 4.5 Solar-Cycle Sensitivity: $B_{r}(1\,\mathrm{AU})$ and $V_{\rm sw}$
4.6 Numerical Stability and Asymptotics
(a) Fractional numerical error
(b) Mean deflection versus outer boundary $r_{\rm stop}$
The numerical tests in Fig. 9(a) quantify the robustness of our trajectory integration scheme and complement the model–comparison figures discussed above. In the left panel, we focus on the Parker unipolar field and track, for several fixed energies, and show how the sky–averaged deflection approaches an asymptotic value as the integration step budget is increased. Expressing the result as the fractional offset $(D-D_{\infty})/D_{\infty}$ relative to the highest–resolution run makes the convergence behavior explicit: for each energy the curves fall monotonically toward zero, and for ${\rm steps/gyro}\gtrsim 1600$ the residual error is below the $1.5\%$ band (and typically at the $\sim 1\%$ level or better). This demonstrates that our production choice (which is more conservative than the minimum satisfying these criteria) yields deflections that are numerically converged for the purposes of our comparison across magnetic geometries.
@@ 5 Discussion
$\langle\Delta\theta\rangle\;\propto\;\frac{\langle|B_{\perp}|\rangle\,L_{\rm eff}}{E}\,$,
so $E_{\rm crit}$ scales linearly with the effective spiral strength and the geometric exposure to sheet encounters. Parameters that raise $E_{\rm crit}$ include larger $\langle|B_{r}(1\,\mathrm{AU})|\rangle$, slower equatorial wind (stronger azimuthal winding), Smith–Bieber–like enhancement of $B_{\phi}$, and HCS configurations that increase encounter probability (larger tilt/waviness). Conversely, configurations that redistribute open flux and weaken the equatorial spiral (e.g., DQCS cores or faster winds toward the poles) lower $E_{\rm crit}$. In our scans, once $E\gtrsim$ a few TeV, all models converge to sub–degree bending for standard normalizations; between $\sim$0.1 and 1 TeV, model choice can shift $\langle\Delta\theta\rangle$ by factors of a few.
@@
Our stability scans (Figs. 9(a) and 9(b)) show that the sky–averaged deflection is well behaved and approaches a clear asymptotic value as the integrator resolution is increased. For the Parker reference case, the fractional deviation $(D-D_{\infty})/D_{\infty}$ falls below $\sim 1$$1.5\%$ once steps_per_gyr exceeds $\simeq 1.5\times 10^{3}$, nearly independent of energy, and continues to decrease for higher step budgets. Independently, varying the outer boundary in the structured DQCS++spiral, wavy-HCS configuration shows that the mean deflection saturates rapidly for $r_{\rm stop}\gtrsim 50$ AU, with changes $\lesssim 0.3\%$ between 50 and 100 AU. Physically, this reflects that most bending is accumulated within the inner $\sim 10$–30 AU, where the large-scale field is strongest and the spiral curvature largest; further out, the field is weaker and predominantly azimuthal, adding little to the net angle. The production settings adopted in this work (step size and $r_{\rm stop}=50$ AU) are chosen to lie safely within these convergence regimes, so the comparative trends between magnetic configurations are controlled by physics rather than by numerical artifacts.

### v2

@@ 4.4 HCS Geometry: Tilt and Waviness
Comparing panels 6(d) and 6(c) shows that replacing a flat sector boundary with a tilted, wavy sheet broadens the 16–84% bands—especially at $E\!\sim\!30$$300$ GeV—while leaving the mean only modestly affected. Physically, additional undulations increase the chance and multiplicity of sector crossings and drift-rich path segments; small changes in launch direction translate into larger path-length and curvature differences across the sky.
@@
Strengthening the local azimuthal field (panel 6(e)) modestly raises sub-TeV means relative to the flat-HCS DQCS+spiral case, consistent with tighter winding, but the difference is again minor compared to the Parker vs. DQCS split from Sec. 4.3. At TeV energies the mean curves nearly coincide.
@@
Introducing a second wave in longitude (panel 6(f)) further amplifies the percentile width at low energies while keeping the mean close to the $m=1$ case. The mean-overlay in Fig. 7 highlights that ordering: Parker remains largest, DQCS(core) is the most suppressed, and the DQCS+spiral variants (flat/wavy, SB) cluster together in mean but differ in spread.
@@ 4.5 Solar-Cycle Sensitivity: $B_{r}(1\,\mathrm{AU})$ and $V_{\rm sw}$
For both the modified Parker model and the configuration including the heliospheric current sheet, the dependence of the mean deflection on the field intensity at 1 AU is qualitatively similar to that of the pure Parker field: $\langle\Delta\theta\rangle\propto B_{1{\rm AU}}^{-1}$ in the ballistic regime and approaches saturation at lower energies. Quantitatively, the presence of latitudinal structure or current-sheet reversals produces at most $\sim$10–15% deviations from this scaling across the 1–10 nT range tested. These variations confirm that our results are robust against moderate uncertainties in the field normalization.
@@ 5 Discussion
$\langle\Delta\theta\rangle\;\propto\;\frac{\langle|B_{\perp}|\rangle\,L_{\rm eff}}{E}\,$, (14)
so $E_{\rm crit}$ scales linearly with the effective spiral strength and the geometric exposure to sheet encounters. The quantity $L_{\rm eff}$ represents the effective magnetic-field path length contributing to the net angular deflection, defined as $L_{\rm eff}\equiv\int(B_{\perp}/B_{1{\rm AU}})\,ds$, where $B_{\perp}$ is the transverse field component along the trajectory. The characteristic energy $E_{\rm crit}$ marks the transition between the strong- and weak-deflection regimes and is approximately given by the condition $r_{\rm L}(E_{\rm crit})\simeq L_{\rm eff}$. For a typical Parker-field normalization, $E_{\rm crit}\!\sim\!200$–300 GeV. Parameters that raise $E_{\rm crit}$ include larger $\langle|B_{r}(1\,\mathrm{AU})|\rangle$, slower equatorial wind (stronger azimuthal winding), Smith–Bieber–like enhancement of $B_{\phi}$, and HCS configurations that increase encounter probability (larger tilt/waviness). Conversely, configurations that redistribute open flux and weaken the equatorial spiral (e.g., DQCS cores or faster winds toward the poles) lower $E_{\rm crit}$. In our scans, once $E\gtrsim$ a few TeV, all models converge to sub–degree bending for standard normalizations; between $\sim$0.1 and 1 TeV, model choice can shift $\langle\Delta\theta\rangle$ by factors of a few.
@@
Our stability scans (Figs. 1(a) and 1(b)) show that the sky–averaged deflection is well behaved and approaches a clear asymptotic value as the integrator resolution is increased. For the Parker reference case, the fractional deviation $(D-D_{\infty})/D_{\infty}$ falls below $\sim 1$$1.5\%$ once steps_per_gyr exceeds $\simeq 1.5\times 10^{3}$, nearly independent of energy, and continues to decrease for higher step budgets. Independently, varying the outer boundary in the structured DQCS++spiral, wavy-HCS configuration shows that the mean deflection saturates rapidly for $r_{\rm stop}\gtrsim 50$ AU, with changes $\lesssim 0.3\%$ between 50 and 100 AU. Physically, this reflects that most bending is accumulated within the inner $\sim 10$–30 AU, where the large-scale field is strongest and the spiral curvature largest; further out, the field is weaker and predominantly azimuthal, adding little to the net angle. The production settings adopted in this work (step size and $r_{\rm stop}=50$ AU) are chosen to lie safely within these convergence regimes, so the comparative trends between magnetic configurations are controlled by physics rather than by numerical artifacts.
@@
We note that for instruments with limited fields of view such as AMS-02 and PAMELA, heliospheric deflections below $\sim$100 GeV are likely to be entangled with magnetospheric effects, which dominate the final few Earth radii of propagation. Dedicated back-tracing including the geomagnetic field would therefore be required to connect observed directions to the heliospheric boundary for sub-TeV cosmic-ray electrons and positrons. Above a few hundred GeV, however, magnetospheric bending becomes negligible ($0.1^{\circ}$), and the present heliospheric modeling suffices to interpret large-scale anisotropy or Sun-shadow observations. We thus recommend combined heliospheric–magnetospheric back-tracing for future low-energy analyses.
@@
We finally also note that at sub–100 GeV energies, the neglected processes of diffusion, drift, and convection become progressively more relevant, potentially broadening and isotropizing the apparent deflection patterns discussed here. Our results should therefore be regarded as upper bounds on the geometric deflections expected in the high-energy, ballistic regime.
@@ A.9 Default Parameters and Scan Ranges
A.10 Code and Reproducibility
  • cre_compare_configs_demo.py (default vs. Parker/DQCS variants),